If f is continuous and "\intop f (x) dx=6 (upper limit=2, lower limit=0)" ,evaluate

"\intop f(2sin\theta) cos\theta d\theta (upper limit=\pi/2 ,lower limit=0)"


1
Expert's answer
2021-09-07T19:00:15-0400

Solution:

Given,"\int_0^2 f(x)dx=6" ...(i)

Let "I=\int_0^{\pi/2} f(2\sin \theta)\cos \theta d\theta"

Put "2\sin \theta=t"

"\Rightarrow 2\cos \theta=\dfrac{dt}{d\theta}"

When "\theta=0,t=0; \theta=\pi/2,t=2"

So, "I=\frac12\int_0^{2}f(t)dt"

"=\frac12\times 6" [Using (i)]

"=3"